# GeoPT: Scaling Physics Simulation via Lifted Geometric Pre-Training

### v1

@@ 2.1 Neural Simulators
Our review focuses on the evolution of neural simulators, which we categorize into model architectures and foundation models. For classical numerical solvers, we refer readers to established reviews (Ŝolín, 2005; Jasak, 2009).
@@
Physics foundation models. Scaling neural simulators as physics foundation models has been explored to improve simulation performance (Yang et al., 2023; Herde et al., 2024; Ye et al., 2024). Poseidon pre-trains on 2D fluid dynamics with temporal conditioning (Herde et al., 2024); DPOT expands to diverse physics with auto-regressive prediction (Hao et al., 2024); Unisolver incorporates PDE information via conditional architectures (Zhou et al., 2025); and P3D extends to 3D fluids (Holzschuh et al., 2025). Despite these advances, existing models remain restricted to a specific physics family on regular grids and do not generalize to the industrial simulations evaluated in this paper. Moreover, they still rely on compute-heavy simulation data for scaling, whereas GeoPT pre-trains on off-the-shelf geometries alone, offering a more scalable path to large-scale pre-training.
@@ 2.2 Self-Supervised Pre-Training
Well-established methods include reconstructing masked inputs (Vincent et al., 2008; He et al., 2022; Xie et al., 2022) and predicting similarity among augmentations (He et al., 2020; Chen et al., 2020; Caron et al., 2021). Despite the diverse pretext tasks, these methods limit the learning process within the pre-training data space without introducing external information. This paradigm is widely adopted for input-aligned pre-training and fine-tuning tasks. Our setting differs fundamentally: we pre-train on geometry yet target generalization to the higher-dimensional physics space, where native-space pre-training may collapse due to potential randomness of uncovered factors, e.g., dynamics.
Self-supervised pre-training has also been explored in 3D geometry understanding (Yu et al., 2022; Tencent, 2025), and recent work incorporates such pre-trained encoders as auxiliary feature extractors for physics-learning (Deng et al., 2024; Zhang et al., 2025). However, these approaches rely on frozen geometric encoders and do not scale the core physics backbone. Furthermore, since pre-training uses static geometric supervision only, the learned representations lack awareness of dynamics. In contrast, we directly pre-train the physics backbone itself and bridge the geometry-physics gap through dynamics-lifted supervision.
@@ 4.1 Lifting Geometry to Physics
Figure 3: Geometry-physics analysis. (a) Visualization of learned correlations on DrivAerML (Ashton et al., 2024). We train Transolver (Wu et al., 2024) using different supervisions and visualize the spatial distribution of learned aggregation weights in four tokens. Brighter colors indicate higher token assignment likelihood, revealing correlations captured by the model. See Appendix G full results. (b) We lift the geometry space by augmenting it with synthetic velocity fields, which further derive a dynamics-aware supervision.
@@
The trajectory evolution in Eq. (4) corresponds to solving a fundamental mass-conservation system, described by the transport equation with sticking boundary: $\partial_{t}f+v\cdot\nabla_{x}f=0$, where $f(x,v,t)$ is a phase-space density. Pre-training with randomly sampled velocity fields $v\in\mathcal{V}$ can be viewed as learning to obey the conservation law under arbitrary dynamics, providing a universal prior for the downstream physical simulation. See Appendix B for details.
@@ 4.3 Implementation Details
We provide key implementation details here; full configurations can be found in Appendix F.
@@
Figure 5: Performance comparison across fine-tuning epochs and physics samples. We show detailed curves at 200 epochs and 100 samples for clarity. Here, geometry-only pre-training adopts vector distance, which is better than SDF. See Appendix D for full results.
5 Experiments
Baselines. GeoPT mainly focuses on the self-supervised pre-training of neural simulators, which has not been well studied previously. Therefore, we adopt native geometry space pre-training: given position information to predict SDF and vector distance, as baselines. Experimentally, since SDF-based pre-training is much worse than vector distance, we defer the SDF-related experiments to Appendix C. Besides, we also compare with the geometry-conditioned paradigm, where we adopt the VAE encoder from the advanced 3D geometry model Hunyuan3D (Tencent, 2025) to extract geometry representations as an auxiliary feature. Additionally, GeoPT is built upon the advanced geometry-general backbone Transolver (Wu et al., 2024). Other Transformer-based simulator backbones, including Galerkin Transformer (2021), GNOT (2023), UPT (2024) and Transolver++ (2025), are also compared.
@@ 5.1 Main Results
Figure 6: GeoPT scaling tests. (a) Gradually increase model layers from 8 to 32 and record the performance change of training from scratch and with GeoPT. (b) Reduce the pre-training diversity of both geometries and dynamics. See Appendix G for full results.
Figure 7: Simulation results and learned representations from GeoPT, including (a) visualization of the prediction results with the worst relative L2 performance in DrivAerML, (b) the error map of surface pressure and surrounding velocity, (c) correlations learned by pre-trained GeoPT under varied dynamics information, such as different directions and speeds of $V_{S}$. See Appendix E for more results.
@@ 5.2 Model Analysis
Generalize to other physics domains. GeoPT is pre-trained with highly diverse dynamics, endowing it with strong potential to generalize across physics domains. We apply GeoPT to radiosity simulation (Goral et al., 1984), a classical light transport problem, which involves fundamentally different governing physics and geometry boundaries from our main experiments. Despite this significant domain shift, GeoPT continues to yield consistent performance improvements. Full results are provided in Appendix A.
@@ 6 Conclusion
Appendix A Extension to Radiosity
To evaluate whether GeoPT transfers to physical regimes beyond those evaluated in our main experiments, we apply it to radiosity simulation (Goral et al., 1984), a classical light transport problem that computes global illumination by modeling diffuse inter-reflections between surfaces. This problem has fundamentally different governing physics from the fluid and solid mechanics tasks in our main experiments. We construct a dataset of 200 radiosity renderings in a Cornell box scene (Goral et al., 1984) with a Stanford bunny placed at varying positions and scales, illuminated by light sources with different intensities, sizes, and directions. We parameterize the dynamics conditions as the light propagation direction, analogous to flow direction in aerodynamics. We finetune GeoPT on 160 training samples and evaluate on 40 held-out test samples. GeoPT achieves an MAE of $9.0\!\times\!10^{-2}$, outperforming training from scratch ($9.7\!\times\!10^{-2}$). Qualitatively, Fig. 9 reveals that GeoPT captures high-frequency shadow boundaries more accurately, particularly in regions with complex light-geometry interactions. Notably, neither the Cornell box geometry nor any light transport physics was seen during pre-training. This result highlights GeoPT’s potential as a general-purpose prior for diverse simulation tasks.
Figure 9: Neural radiosity simulation on Cornell box with Stanford bunny. GeoPT captures high-frequency shadow boundaries (dashed boxes) more accurately than training from scratch.
Appendix B Theoretical Understanding of GeoPT Pre-Training
@@
It can be proved that, Eq. (9) describes advection in phase space along characteristic curves (Evans, 2022)
@@
Proposition B.1 (Mass conservation).
@@
Remark B.2 (Connect between GeoPT pre-training and solving collisionless transport equation in Eq. (9)).
@@
As stated in proposition B.1, the flow density $f$ that is used as a self-supervision signal obeys mass conservation under all kinds of velocity settings. During the pre-training process, we randomly sample geometries $g$ and velocity fields $\mathbf{v}$ as the model input. Beyond ensuring data diversity, the above-described pre-training will enforce the model output maintain the mass conservation law under all kinds of geometries and dynamics.
Remark B.3 (Generalizability of theoretical framework).
@@
Appendix C Pre-Training Investigation
@@
Appendix D Fine-Tuning Investigation
@@
Remark D.1 (Distinguishable configurations).
@@
Appendix E Showcases
As a supplement to Fig. 7, we plot the showcases from the other four benchmarks in Fig. 15-18.
@@ Appendix E Showcases
Appendix F Implementation Details
@@
F.1 Benchmarks
We experiment with five industrial design tasks, where DrivAerML (Ashton et al., 2024), NASA-CRM (Bekemeyer et al., 2025) and AirCraft (Luo et al., 2025) are adopted from the previous work. We also newly simulate two benchmarks.
@@ F.1 Benchmarks
F.2 Self-Supervision Data
Algorithm 1 Self-supervision data generation in GeoPT
Input Data: Geometry dataset $\mathcal{G}$
 Initialize dataset $\mathcal{D}\leftarrow\emptyset$
for $G$ in $\mathcal{G}$ do
  // Step 1: Normalize geometry
  // Step 2: Sample query positions
  for $i$ in $\{1,\ldots,N_{\text{dyn}}\}$ do
   // Step 3: Sample synthetic velocity field
   // Step 4: Compute feature trajectory following Eq. (4) and Eq. (5)
   $\{\mathbf{x}_{0}\}\leftarrow\{\mathbf{x}\}$
   for $t$ in $\{0,\ldots,\tau\}$ do
   end for
  end for
end for
F.3 Experiment Configuration
@@
F.4 Baselines
@@
Since prior research does not directly study geometric pre-training for physics, we compare GeoPT with two constructed baselines. Here are details for these two geometry usage baselines.
@@
(ii) Geometry-only conditioning. The prior work (Zhang et al., 2025) has explored using frozen geometry representation as the auxiliary information for PDE solving. However, the 3D model used in their work is limited to 3D inductors, which cannot be used in complex industrial designs tested in our paper. Thus, we adopt the advanced geometry model Hunyuan3D (Tencent, 2025) to extract the static geometry representation for comparison. Specifically, we adopt the pre-trained VAE model [^4] https://huggingface.co/tencent/Hunyuan3D-2/tree/main/hunyuan3d-vae-v2-0-withencoder encoder, which can receive a set of points sampled from a mesh geometry as input and encode it into 3,072 geometry tokens with 64 hidden channels. Then we integrate the extracted token sequence into Transolver based on an additional cross-attention layer to fuse the geometry tokens into input mesh representations, which is also the default usage in (Zhang et al., 2025). As shown in Fig. 20, the geometry representation learned by Hunyuan3D can precisely capture the detailed geometry information, thereby enabling accurate geometry reconstruction. However, as presented in the main results (Fig. 5), the static geometry representation cannot help the physics learning, even though it has already precisely encoded the geometry information.
Appendix G Full Results

### v2

@@ 2.1 Neural Simulators
Our review focuses on the evolution of neural simulators, which we categorize into model architectures and foundation models. For classical numerical solvers, we refer readers to established reviews (Solin, 2005; Jasak, 2009).
@@
Physics foundation models. Scaling neural simulators as physics foundation models has been explored to improve simulation performance (Yang et al., 2023; Herde et al., 2024; Ye et al., 2024). Poseidon pre-trains on 2D fluid dynamics with temporal conditioning (Herde et al., 2024); DPOT expands to diverse physics with auto-regressive prediction (Hao et al., 2024); Unisolver incorporates PDE information via conditional architectures (Zhou et al., 2025); and P3D extends to 3D fluids (Holzschuh et al., 2026). Despite these advances, existing models remain restricted to a specific physics family on regular grids and do not generalize to the industrial simulations evaluated in this paper. Moreover, they still rely on compute-heavy simulation data for scaling, whereas GeoPT pre-trains on off-the-shelf geometries alone, offering a more scalable path to large-scale pre-training.
@@ 2.2 Self-Supervised Pre-Training
Well-established methods include reconstructing masked inputs (Vincent et al., 2008; He et al., 2022; Xie et al., 2022) and predicting similarity among augmentations (He et al., 2020; Chen et al., 2020; Caron et al., 2021). Despite the diverse pretext tasks, these methods limit the learning process within the pre-training data space without introducing external information. This paradigm is widely adopted for input-aligned pre-training and fine-tuning tasks. Our setting differs fundamentally: we pre-train on geometry yet target generalization to the higher-dimensional physics space, where native-space pre-training may collapse due to potential randomness of uncovered factor, e.g., dynamics.
Self-supervised pre-training has also been explored in 3D geometry understanding (Yu et al., 2022; Tencent, 2025), and recent work incorporates such pre-trained encoders as auxiliary feature extractors for physics-learning (Deng et al., 2024; Zhang et al., 2026). However, these approaches rely on frozen geometric encoders and do not scale the core physics backbone. Furthermore, since pre-training uses static geometric supervision only, the learned representations lack awareness of dynamics. In contrast, we directly pre-train the physics backbone itself and bridge the geometry-physics gap through dynamics-lifted supervision.
@@ 4.1 Lifting Geometry to Physics
Figure 3: Geometry-physics analysis. (a) Visualization of learned correlations on DrivAerML (Ashton et al., 2024). We train Transolver (Wu et al., 2024) using different supervisions and visualize the spatial distribution of learned aggregation weights in four tokens. Brighter colors indicate higher token assignment likelihood, revealing correlations captured by the model. See Appendix F full results. (b) We lift the geometry space by augmenting it with synthetic velocity fields, which further derive a dynamics-aware supervision.
@@
The trajectory evolution in Eq. (4) corresponds to solving a fundamental mass-conservation system, described by the transport equation with sticking boundary: $\partial_{t}f+v\cdot\nabla_{x}f=0$, where $f(x,v,t)$ is a phase-space density. Pre-training with randomly sampled velocity fields $v\in\mathcal{V}$ can be viewed as learning to obey the conservation law under arbitrary dynamics, providing a universal prior for the downstream physical simulation. See Appendix A for details.
@@ 4.2 Lifted Geometric Pre-Training
Remark 4.2 (Native space v.s. lifted pre-training).
By comparing native space (Eq. (2)) and lifted (Eq. (6)) pre-training formalizations, we can establish the theoretical connection between these two paradigms. Given optimal models $\mathcal{F}_{{\theta}_{\text{native}}^{\ast}}$ and $\mathcal{F}_{{\theta}_{\text{lifted}}^{\ast}}$ achieving zero $\mathcal{L}^{\text{pre}}_{\text{native}},\mathcal{L}^{\text{pre}}_{\text{lifted}}$ respectively, we have
$\text{slice}\left(\mathbb{E}_{V}\left[\mathcal{F}_{{\theta}_{\text{lifted}}^{\ast}}(\mathbf{x};G,V)\right]\right)=\mathcal{F}_{{\theta}_{\text{native}}^{\ast}}(\mathbf{x};G),$ (8)
where slice denotes taking $t=0$. This can be easily proven based on the definition of $\boldsymbol{h}_{G}$ in Eq. (5). This equation indicates that the canonical native-space pretraining essentially learns the expectation of our lifted paradigm, which may lead to information loss and degenerated representations. In contrast, lifted pre-training learns the spanned manifold, where the dynamics $V$ can also serve as the “key” to retrieve corresponding capability learned during pre-training.
@@ 4.3 Implementation Details
We provide key implementation details here; full configurations can be found in Appendix E.
@@
Figure 5: Performance comparison across fine-tuning epochs and physics samples. We show detailed curves at 200 epochs and 100 samples for clarity. Here, geometry-only pre-training adopts vector distance, which is better than SDF. See Appendix C for full results.
5 Experiments
Baselines. GeoPT mainly focuses on the self-supervised pre-training of neural simulators, which has not been well studied previously. Therefore, we adopt native geometry space pre-training: given position information to predict SDF and vector distance, as baselines. Experimentally, since SDF-based pre-training is much worse than vector distance, we defer the SDF-related experiments to Appendix B. Besides, we also compare with the geometry-conditioned paradigm, where we adopt the VAE encoder from the advanced 3D geometry model Hunyuan3D (Tencent, 2025) to extract geometry representations as an auxiliary feature. Additionally, GeoPT is built upon the advanced geometry-general backbone Transolver (Wu et al., 2024). Other Transformer-based simulator backbones, including Galerkin Transformer (2021), GNOT (2023), UPT (2024) and Transolver++ (2025), are also compared. All methods are under the same training strategy to ensure a fair comparison.
@@ 5.1 Main Results
Figure 6: GeoPT scaling tests. (a) Gradually increase model layers from 8 to 32 and record the performance change of training from scratch and with GeoPT. (b) Reduce the pre-training diversity of both geometries and dynamics. See Appendix F for full results.
Figure 7: Simulation results and learned representations from GeoPT, including (a) visualization of the prediction results with the worst relative L2 performance in DrivAerML, (b) the error map of surface pressure and surrounding velocity, (c) correlations learned by pre-trained GeoPT under varied dynamics information, such as different directions and speeds of $V_{S}$. See Appendix D for more results.
@@ 5.2 Model Analysis
Figure 9: Experiments of neural radiosity simulation on the Cornell box (Goral et al., 1984) with an unseen Stanford bunny geometry.
@@
5.3 Towards Physics Foundation Model
Generalize to other physics domains. GeoPT is pre-trained with highly diverse dynamics, endowing it with strong potential to generalize across physics domains. We apply GeoPT to radiosity simulation (Goral et al., 1984), a classical light transport problem, which involves fundamentally different governing physics and geometries from our main experiments. Despite a significant domain shift, GeoPT continues to yield consistent performance improvements, which achieves an MAE of $9.0\!\times\!10^{-2}$, outperforming training from scratch ($9.7\!\times\!10^{-2}$). Qualitatively, Fig. 9 reveals that GeoPT captures high-frequency shadow boundaries more accurately, particularly in regions with complex light-geometry interactions. More implementation details of this task can be found in Appendix E.1.
Pre-training on general-purpose geometries. As shown in Fig. 6, GeoPT exhibits favorable scalability, suggesting its potential to be further developed into a foundation model (Bommasani et al., 2021) for physics simulation. To further investigate this potential, we enhance GeoPT by increasing the diversity of pre-training geometries. Specifically, we pre-train GeoPT under the following three settings:
(i) ShapeNet subsets ($\sim$10,000 geometries). We use the car, airplane, and watercraft subsets from ShapeNet, which is also the default pre-training setting in previous experiments.
(ii) ShapeNet full ($\sim$50,000 geometries). As a general-purpose 3D shape repository, ShapeNet (2015) contains 55 object categories, covering a wide range of shapes such as mobile devices and furniture. Here, we use the full dataset.
(iii) ShapeNet+Hunyuan3D ($\sim$300,000 geometries). Beyond public 3D assets, advanced 3D generative models can provide an additional source of diverse pre-training geometries. We therefore augment the pre-training data with HY3D-Bench (Hunyuan3D et al., 2026), which contains over 250k geometries generated by Hunyuan3D (2025).
As shown in Fig. 10, GeoPT can be further improved by increasing the diversity of pre-training geometries, even when these geometries differ substantially from those used during fine-tuning, especially for AirCraft with rarely observed smooth geometries. This result suggests that GeoPT can effectively benefit from large-scale and diverse geometries, providing a practical path toward scaling up pre-training.
Together, the favorable physics and geometry generalizability demonstrated above highlight the potential of GeoPT as a step toward general-purpose physics foundation models.
Figure 10: Scaling up GeoPT-Huge pre-training with general 3D shape repositories. Left: performance under different pre-training scales. Right: geometry examples from the pre-training datasets.
@@ 6 Conclusion
Acknowledgements
We would like to thank Hang Zhou and Yuezhou Ma from Tsinghua University for their helpful discussion during the initial exploration of this project.
@@ Acknowledgements
Appendix A Theoretical Understanding of GeoPT Pre-Training
@@
It can be proved that, Eq. (10) describes advection in phase space along characteristic curves (Evans, 2010)
@@
Proposition A.1 (Mass conservation).
@@
Remark A.2 (Connect between GeoPT pre-training and solving collisionless transport equation in Eq. (10)).
@@
As stated in proposition A.1, the flow density $f$ that is used as a self-supervision signal obeys mass conservation under all kinds of velocity settings. During the pre-training process, we randomly sample geometries $g$ and velocity fields $\mathbf{v}$ as the model input. Beyond ensuring data diversity, the above-described pre-training will enforce the model output maintain the mass conservation law under all kinds of geometries and dynamics.
Remark A.3 (Generalizability of theoretical framework).
@@
Appendix B Pre-Training Investigation
@@
Appendix C Fine-Tuning Investigation
@@
Remark C.1 (Distinguishable configurations).
@@
Appendix D Showcases
As a supplement to Fig. 7, we plot the showcases from the other four benchmarks in Fig. 16-19.
@@ Appendix D Showcases
Appendix E Implementation Details
@@
E.1 Benchmarks
We experiment with five industrial design tasks and one radiosity task to examine the model generalizability, where DrivAerML (Ashton et al., 2024), NASA-CRM (Bekemeyer et al., 2025) and AirCraft (Luo et al., 2025) are adopted from the previous work. We also newly simulate three benchmarks, including DTCHull, Car-Crash and Radiosity.
@@ E.1 Benchmarks
Radiosity
To evaluate whether GeoPT transfers to physical regimes beyond those evaluated in our main experiments, we apply it to radiosity simulation (Goral et al., 1984), a classical light transport problem that computes global illumination by modeling diffuse inter-reflections between surfaces. This problem has fundamentally different governing physics from the fluid and solid mechanics tasks in our main experiments. We construct a dataset of 200 radiosity renderings in a Cornell box scene (Goral et al., 1984) with a Stanford bunny placed at varying positions and scales, illuminated by light sources with different intensities, sizes, and directions. We parameterize the dynamics conditions as the light propagation direction, analogous to the flow direction in aerodynamics. We finetune GeoPT on 160 training samples and evaluate on 40 held-out test samples. Notably, neither the Cornell box geometry nor any light transport physics was seen during pre-training. This benchmark examines GeoPT’s potential as a general-purpose prior for diverse simulation tasks.
E.2 Experiment Configuration
@@
E.3 Self-Supervision Data
Algorithm 1 Self-supervision data generation in GeoPT
Input Data: Geometry dataset $\mathcal{G}$
 Initialize dataset $\mathcal{D}\leftarrow\emptyset$
for $G$ in $\mathcal{G}$ do
  // Step 1: Normalize geometry
  // Step 2: Sample query positions
  for $i$ in $\{1,\ldots,N_{\text{dyn}}\}$ do
   // Step 3: Sample synthetic velocity field
   // Step 4: Compute feature trajectory following Eq. (4) and Eq. (5)
   $\{\mathbf{x}_{0}\}\leftarrow\{\mathbf{x}\}$
   for $t$ in $\{0,\ldots,\tau\}$ do
   end for
  end for
end for
E.4 Baselines
@@
Since prior research does not directly study geometric pre-training for physics, we compare GeoPT with two reasonable baselines. Here are details for these two geometry usage baselines.
@@
(ii) Geometry-only conditioning. The prior work (Zhang et al., 2026) has explored using frozen geometry representation as the auxiliary information for PDE solving. However, the 3D model used in their work is limited to 3D inductors, which cannot be used in complex industrial designs tested in our paper. Thus, we adopt the advanced geometry model Hunyuan3D (Tencent, 2025) to extract the static geometry representation for comparison. Specifically, we adopt the pre-trained VAE model [^4] https://huggingface.co/tencent/Hunyuan3D-2/tree/main/hunyuan3d-vae-v2-0-withencoder encoder, which can receive a set of points sampled from a mesh geometry as input and encode it into 3,072 geometry tokens with 64 hidden channels. Then we integrate the extracted token sequence into Transolver based on an additional cross-attention layer to fuse the geometry tokens into input mesh representations, which is also the default usage in (Zhang et al., 2026). As shown in Fig. 21, the geometry representation learned by Hunyuan3D can precisely capture the detailed geometry information, thereby enabling accurate geometry reconstruction. However, as presented in the main results (Fig. 5), the static geometry representation cannot help the physics learning, even though it has already precisely encoded the geometry information.
Appendix F Full Results
Extension to time-dependent simulation
[…]
Extension to other backbones
As demonstrated in Figure 8(b), we choose Transolver as the default backbone of GeoPT, whose performance is also verified in recent research (Zhou et al., 2026). It is worth noticing that GeoPT is an architecture-agnostic pre-training model. Therefore, it is also possible to apply lifted geometric pre-training to GNOT (Hao et al., 2023) and Galerkin Transformer (Cao, 2021). As shown below, GeoPT can also bring improvements to these models.
Table 4: GeoPT Extension to time-dependent simulation and other backbones. All metrics are tested under full-data and full-training.
(a) Time-dependent simulation of DTCHull.
Relative L2 Transient DTCHull
From scratch 0.06189
GeoPT 0.05402
(b) Applying GeoPT to other backbones.
Relative L2 DrivAerML NASA-CRM AirCraft DTCHull Car-Crash
Galerkin (2021) 0.127 0.109 0.136 0.210 0.208
GeoPT-Galerkin 0.083 0.095 0.099 0.156 0.182
GNOT (2023) 0.152 0.105 0.122 0.211 0.400
GeoPT-GNOT 0.116 0.097 0.099 0.180 0.219